In problem R=√R²×+R²y=√(3.0km)²+(1.0km)²

R=2.24km, in what direction she from the starting point?


1
Expert's answer
2021-02-05T11:47:39-0500

If


"R=\sqrt{R^2_x+R^2_y}=\sqrt{(3.0\text{ km})²+(1.0\text{ km})²},"

then


"R=\sqrt{10}=3.16\text{ km}."

This vector (R) makes the following angle with the 3km-vector:


"\theta=\text{atan}\frac{1}{3}=18.4°."